OpenStudy (anonymous):

Let -pi/4 < x < pi/4 , f(x)=tan(2x). A tangent to the graph of y=f(x) where x= a makes an angle of 70 degrees with the positive direction of the x-axis. Find the possible values of a

OpenStudy (anonymous):

x=a=70 degrees = 7pi/18

OpenStudy (rational):

HINT : slope = f'(a) = tan(70)

OpenStudy (anonymous):

could you please explain it more i know that the gradient is equal to f'(a) but what i couldnt understand was tan(70)

OpenStudy (rational):

whats the definition of "tan" ? remember SOH-CAH-TOA ?

OpenStudy (anonymous):

opposite/adjacent

OpenStudy (rational):



OpenStudy (rational):

and whats the definition of slope ?

OpenStudy (rational):

how do u get slope of a line ?

OpenStudy (anonymous):

derivative

OpenStudy (rational):

suppose you don't know calculus, what would u do to get the slope of a line ?

OpenStudy (rational):

slope = (y2-y1) / (x2 - x1) right ?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

right

OpenStudy (rational):